HUBBLE SPACE TELESCOPE DAILY REPORT #2243 PERIOD COVERED: 0000Z (UTC) 11/03/98 - 0000Z (UTC) 11/04/98 Daily Status Report as of 308/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Twelve Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.3 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.4 Completed STIS/CCD 7354 (Kinematics of Gaseous Disks in the Nuclei of Liners with Compact Flat Spectrum Radio Sources) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclei of the liner in the compact flat spectrum source galaxy NGC-6500. Many active galactic nuclei channel energetic particles into jets that feed powerful extended lobes of radio emission; others remain quite compact. Both classes of active galaxies are suspected to be powered gravitationally by the accretion of surrounding material onto a massive black hole. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the star HP-TAUG2 as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS (CCD/MA1/MA2) 7346 (The Nature of the Bright 'Supersodt X-ray Source' RX J 0019+21) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to obtain decisive information on the true nature of RX J 0019+21, a bright "supersoft x-ray source". Two sets of echelle spectra at the quadrature phases of the binary orbit in order to determine the UV spectral energy distribution, to separate stellar and interstellar components, to search for information on the absorption of stellar or disk origin, to measure LineHell1160 emission lines, and to search for other emission/absorption lines. The proposal completed with no reported problems. 1.8 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell V992-SCO in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.9 Completed NIC/2 7863 (The Nature of Gamma-Ray Burster Host Galaxies) The Near Infrared Camera (Camera Two) was used to study the nature of a gamma-ray burster in the host galaxy GRB970508. Of the two thousand gamma-ray bursts detected in the last twenty-five years, only two have identified optical counterparts. The observations were executed as scheduled, and no problems were reported. 1.10 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.11 Completed NIC/1/2/3 7226 (Search for Massive Jupiters) The Near Infrared Camera (Cameras One, Two, and Three) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/1/2/3 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Cameras One, Two, and Three) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.13 Completed FGS 7678 (Monitor the Long-Term Stability of FGS-1R) The goal of these measurements is to monitor the internal positional stability of the FGSs by measuring many stars in the M35 check field repeatedly over a one year period. The results of these measurements are the stability characteristics of the FGS to allow an estimate of differential corrections to OFAD and Plate Scale. The observations were completed as planned, and no problems were reported. 1.14 Completed Three Sets of NIC/1/2/3 7336 (Cosmological Parameters Omega and Lambda From High-Redshift Type Ia Supernovae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high redshifted Type Ia supernova SN98104. HST can make a direct measurement of the cosmological parameters, Omega, Lambda, and curvature, using Type Ia supernovae as calibrated "standard candles." The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.16 Completed Three Sets of NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 7278 (The Connection Between The Obscuring Torus and the Masing Disk in H_20 Me) The WF/PC-2 was used to investigate the connection between the molecular torus and the emission line gas in a complete sample of H_2O megamasers. The proposal completed nominally with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 23 Successful : 23 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared four times. Gyro #2 was powered up at 307/1447Z as directed by an opertions request. Initial telemetry was nominal, indicating good sync and acceptable temperatures. The motor current profile was similar to the last power-on event, initially rising to 140 mA and then gradually settling to 134 mA. The heater duty cycle was approximately 30%. The performance is being monitored closely pending the expected reconfiguration to a four-gyro configuration on day 310 (11/6/98). Following Gyro #2 turn-on, the Gyro #1 motor current shifted down approximately 3 mA to 145 mA for one telemetry sample (six seconds) at 307/1503Z. HSTAR 6755 was written. It is believed that the shift may have been caused by a slight thermal perturbation resulting from the activation of Gyro #2. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The Battery #2 capacity test is scheduled to begin at approximately 308/1835Z. Continuation of normal science observations and calibrations. /DMH